OpenStudy (anonymous):

A bag contains 10 marbles of the same size that are red, yellow, and orange. The probability of picking a red marble is 10%. The probability of picking a yellow marble is 60%. Determine the probability of picking an orange marble from the bag and then classify the probability of picking an orange marble as likely or not likely to occur.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

anyone???

OpenStudy (abdullahm):

There are only 3 marble colors - red, yellow and orange

OpenStudy (michele_laino):

Hint: the sum of your probabilities has to be equal to 1

OpenStudy (abdullahm):

So, to find out how much for orange we have to do 100% - probability for both red and yellow

TheSmartOne (thesmartone):

Or, if you want to simplify the percents to decimal, you can do as Michele said :)

OpenStudy (dan815):

hunnid

OpenStudy (anonymous):

@AbdullahM has given you the hint

OpenStudy (dan815):

hunnid dolla bills

OpenStudy (anonymous):

Okay i think i got it thanks guysss c:

OpenStudy (michele_laino):

hint: we have top solve this equation: \[x + 60 + 10 = 100\] where x is the requested probability

OpenStudy (abdullahm):

Did you get an answer? We can check it for you :)

rvc (rvc):

All the best :)

OpenStudy (michele_laino):

oops..we have to solve...

OpenStudy (anonymous):

if the total marbles(10) = 100%, then 1% = 10/100 = 0.1 and 10% are red marbles = 10 x 0.1 and 60% are yellow= 60 x 0.1 then that leaves 30% for the yellow marbles

OpenStudy (anonymous):

sorry for the orange marbles (30%)

OpenStudy (anonymous):

P(red marble) = 1/10 x 100 = 10% P(yellow marble) = 6/10 x100 = 60% P(orange marble)= 3/10 x100 = 30%

OpenStudy (anonymous):

sorry to jump you @Michele_Laino